Q23 Suppose H0 is left-tailed and tTest = -2.26. (Note tTest not zTest). The tTest value was calculated from a sample of size n = 13. Which statement below is true? Select one:

a. Fail to reject H0 if level of significance = 0.022

b. Reject H0 if level of significance = 0.022

c. Reject H0 if level of significance = 0.03

d. Fail to reject H0 if level of significance = 0.01
Answer #1

Solution :

This is the left tailed test .

n = 13

degrees of freedom = n - 1 = 13 - 1 = 12

Test statistic = -2.26

P-value = 0.0216

= 0.022  

P-value <

Reject the null hypothesis .

b. Reject H0 if level of significance = 0.022